The reaction of bromine gas with chlorine gas has a Kc value of 7.20 at 200°C. If a closed 2.50 L vessel were charged with the two reactants, with the former weighing 9.65 g while the latter is 5.70 g, what would be the equilibrium concentration of the two reactants and BrCl(g)? Determine whether the forward or the reverse reaction will be favored. Explain your answer.


1
Expert's answer
2022-04-11T09:13:04-0400

Solution:

The balanced chemical equation:

Br2(g) + Cl2(g) = 2BrCl(g)

ICE (Initial, Change, Equilibrium) table for this reaction:



[Br2] = [Cl2] = (0.2 - x) M

[BrCl] = (2x) M

The equilibrium constant expression Kc for the balanced equation:



7.20 = [2x / (0.2 - x)]2

2.6833 = 2x / (0.2-x)

2.6833 × (0.2 - x) = 2x

0.53666 - 2.6833x = 2x

0.53666 = 4.6833x

x = 0.1146


[Br2] = [Cl2] = 0.2 - x = 0.2 - 0.1146 = 0.0854 M

[BrCl] = 2x = (2 × 0.1146) = 0.2292 M

[BrCl] = 0.2292 M